OpenStudy (anonymous):

One yard company charges $10x + $50 a month to mow and edge a lawn. A second company charges $8x + $60 a month to mow and edge a lawn. Each lawn company also offers additional services for an extra fee. Let x represent the number of additional services. Mr. Mead is comparing the two companies. How many additional services are needed for the two companies to be the same price? Mr. Mead needs x additional services for the companies to cost the same price.

OpenStudy (jordanloveangel):

set the two equations equal to each other and solve for x 10x +50 = 8x +60 -50 -50 ----------------- 10x = 8x + 10 10x = 8x + 10 -8x -8x --------------- 2x = 10 2x = 10 divide by 2 on both sides x=5 five additional services are needed for the two companies prices to be equal

OpenStudy (anonymous):

thanks.